# M2NO: Multiresolution Operator Learning with Multiwavelet-based Algebraic Multigrid Method

### v2

@@
Abstract
Solving partial differential equations (PDEs) effectively necessitates a multi-scale approach, particularly critical in high-dimensional scenarios characterized by increasing grid points or resolution. Traditional methods often fail to capture the detailed features necessary for accurate modeling, presenting a significant challenge in scientific computing. In response, we introduce the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), a novel deep learning framework that synergistically combines multiwavelet transformations and algebraic multigrid (AMG) techniques. By exploiting the inherent similarities between these two approaches, M2NO overcomes their individual limitations and enhances precision and flexibility across various PDE benchmarks. Employing Multiresolution Analysis (MRA) with high-pass and low-pass filters, the model executes hierarchical decomposition to accurately delineate both global trends and localized details within PDE solutions, supporting adaptive data representation at multiple scales. M2NO also automates node selection and adeptly manages complex boundary conditions through its multiwavelet-based operators. Extensive evaluations on a diverse array of PDE datasets with different boundary conditions confirm M2NO’s superior performance. Furthermore, M2NO excels in handling high-resolution and super-resolution tasks, consistently outperforming competing models and demonstrating robust adaptability in complex computational scenarios.
1 Introduction
Partial differential equations (PDEs) are fundamental to modeling a wide range of phenomena across science and engineering [13], including fluid dynamics [28, 24], material stress analysis [32], and atmospheric conditions. Traditionally, solving these equations requires specific numerical methods, designed to address individual problem characteristics, often resulting in significant computational efforts [17]. Recent developments in deep learning have introduced alternative approaches to solving PDEs [25, 22, 23, 8, 21], featuring methods like Fourier Neural Operators (FNO) [22], Galerkin Transformers (GT) [8] and Deep Operator Networks (DeepONet)[23]. These techniques leverage neural network architectures to approximate the mappings between parameters and solutions in infinite-dimensional spaces, efficiently addressing both forward and inverse problems in PDEs. Among these techniques, the Fourier Neural Operators (FNO) [22, 20] stand out as a particularly significant development. FNO integrates Fast Fourier Transforms within a neural network framework, effectively learning direct mappings between infinite-dimensional function spaces. This capability marks FNO as a powerful tool, offering a scalable and efficient alternative to traditional numerical methods for solving PDEs.
A primary limitation of the Fourier Neural Operator (FNO) is its reliance on basis functions generated by Fast Fourier Transforms (FFTs). These functions are localized in frequency but lack spatial resolution, leading to difficulties in managing complex boundary conditions and capturing local features in high-dimensional PDEs [2, 9]. In contrast, wavelets provide both spatial and frequency localization [9, 2, 3], which enhances the management of discontinuities and improves image pattern recognition over FFTs. Despite these advantages, the Multiwavelet Transformation (MWT) model, which utilizes wavelet kernels computed by four independent neural networks, faces its own set of challenges. While leveraging Multi-Resolution Analysis (MRA) to provide a detailed representation of both frequency and spatial information, MWT requires substantial computational resources during training and struggles with grid sizes that are not powers of two, which limits its generalizability and practical application [12].
This limitation in MWT serves as a segue into discussing the Algebraic Multigrid (AMG) method. Although AMG is highly effective in improving computational efficiency in numerical computing, it is traditionally constrained by its compatibility with only simple periodic boundary conditions. Moreover, AMG requires the pre-determination of mapping operators between grid points at each level, which poses further limitations on its generalizability and practicality [6, 35, 29, 4].
In response to these limitations, we introduce an innovative neural network architecture, the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), which synergistically combines the strengths of algebraic multigrid methods and multiresolution analysis [5, 11, 7]. Our contributions are designed to overcome the specific challenges faced by previous models:
  • Enhanced Operator Learning: M2NO merges algebraic multigrid and multiresolution techniques to facilitate adaptive learning across multiple resolutions. This integration not only enables automatic node selection but also significantly improves the handling of complex PDE coefficients and intricate boundary conditions.
  • Advanced Operator Design: We develop advanced prolongation and restriction operators based on multiwavelet theory. These operators are critical components of our neural operator framework, enabling efficient hierarchical decomposition that captures both global and local structural details within PDE solutions.
  • Empirical Validation: M2NO is rigorously tested across a variety of 1D and 2D PDE datasets under different boundary conditions. Our experiments confirm its superior accuracy and scalability, consistently outperforming traditional models at various resolution scales.
These innovations underscore our commitment to advancing PDE computational frameworks, offering robust solutions for complex mathematical challenges in scientific computing.
2 Algebraic Multigrid Method and Multiwavelet
@@
In this section, we first provide a concise overview of the algebraic multigrid method [35, 6] and multiresolution analysis using multiwavelets [1, 9]. We then summarize the similarities between these approaches and illustrate how they inform our development of multiwavelet-based prolongation and restriction operators.
2.1 Algebraic Multigrid Method
@@
The algebraic multigrid method is a well-grounded numerical technique for solving systems of algebraic equations that approximate partial differential equations (PDEs). Specifically, we address the problem of solving the linear equation system:
$A^{h}u^{h}=b^{h},$ (1)
@@
$u^{h}=u^{h}+I_{2h}^{h}e^{2h}$ (3)
and perform another few relaxation steps (typically one or two) on the fine grid. This procedure outlines the foundational steps of the V-cycle multigrid scheme, which is recursively defined, as shown in Algorithm 1. For a comprehensive discussion and theoretical grounding of the Algebraic Multigrid (AMG) methods, please refer to the Appendix A.
2.2 Multiwavelets
@@
A multiresolution analysis (MRA) using multiwavelets involves a sequence of closed subspaces, defined as:
@@
$\mathbf{V}_{n}^{k}\oplus\mathbf{W}_{n}^{k}=\mathbf{V}_{n+1}^{k},\quad\mathbf{W}_{n}^{k}\bot\mathbf{V}_{n}^{k},$ (4)
and the dimension of $\mathbf{W}_{n}^{k}$ is of dimension $2^{n}k$. Therefore, the composition of spaces is:
$\mathbf{V}_{n}^{k}=\mathbf{V}_{0}^{k}\oplus\mathbf{W}_{0}^{k}\oplus\mathbf{W}_{1}^{k}\oplus\ldots\oplus\mathbf{W}_{n-1}^{k}.$ (5)
@@
In the two-dimensional discrete wavelet transform, depicted in Fig 4(a), $H$ and $G$ function as low-pass and high-pass filters, respectively, allowing the partitioning of frequency components at each level. Fi 4(b) visually represents this decomposition in a matrix format, enhancing our understanding of how submatrices $B_{n}$, $C_{n}$, and $D_{n}$ are derived from $L_{n}$.
The wavelet transform is defined as $\mathcal{W}:V_{j}\to V_{j-1}\oplus W_{j-1}$ (for $j\in\mathbb{Z}^{+}$), and given by:
$\mathcal{W}=\begin{pmatrix}H\\ G\end{pmatrix}$ (6)
@@
2.3 Similarities between algebraic multigrid and multiresolution
@@
  1. 1.
    The highest resolution space $V_{n}$ in Multiresolution Analysis (MRA) aligns with $\Omega^{h}$, which represents the space of the finest grid vectors in the multigrid method.
  2. 2.
    The decomposition step in MRA, where data is broken down, corresponds to the restriction operation in multigrid methods. Similarly, the forward wavelet transform aligns with the restriction operator $I_{h}^{2h}$.
  3. 3.
    The reconstruction step in MRA, which reassembles the data, is equivalent to the prolongation operation in multigrid methods. Correspondingly, the inverse wavelet transform is equivalent to the prolongation operator $I_{2h}^{h}$.
  4. 4.
    The orthogonality condition $\mathcal{W}\mathcal{W}^{T}=I$ corresponds to the identity relationship for the grid transfer operators, $I_{h}^{2h}I_{2h}^{h}=I$.
  5. 5.
@@
These comparisons highlight the synergy between the algebraic multigrid and multiresolution properties of wavelet transforms and the error smoothing and correcting capabilities of intergrid operators in multigrid methods.
Figure 1: Illustration of the multiwavelet-based multigrid method for PDEs. The diagram outlines the method’s workflow, starting with the initial function $f_{n}$ and guess $u_{n}^{0}$. It includes pre-smoothing, multiwavelet-based restriction, solution of correction equations, prolongation back to finer grids, and post-smoothing, illustrating the data flow and transformations within the multigrid process.
3 Multiwavelet-based Algebraic Multigrid Neural Operator
Building on the synergistic qualities of algebraic multigrid and multiresolution wavelet transforms discussed previously, this section elaborates on the Multiwavelet-based Multigrid Neural Operator (M2NO). We begin by defining the problem and its theoretical basis in section 3.1, then detail the construction and practical implementation of multiwavelet-based operators in section 3.2, and finally, illustrate the diverse applications of these methodologies in solving discrete PDEs in section 3.3.
3.1 Problem definition
@@
Formally, let $\mathcal{A}$ and $\mathcal{U}$ denote two Sobolev spaces $\mathcal{H}^{s,p}$ with $s>0$ and $p\geq 1$. In this context, the operator $T$ maps from $\mathcal{A}$ to $\mathcal{U}$, $T:\mathcal{A}\to\mathcal{U}$. We focus on cases where $s>0$ and $p=2$, choosing $p=2$ specifically to enable the definition of projections within a Hilbert space structure, using measures $\mu$.

### v3

# M2NO: An Efficient Multi-Resolution Operator Framework for Dynamic Multi-Scale PDE Solvers

@@
Abstract.
Solving high-dimensional partial differential equations (PDEs) efficiently requires handling multi-scale features across varying resolutions. To address this challenge, we present the Multiwavelet-based Multigrid Neural Operator (M2NO), a deep learning framework that integrates a multigrid structure with predefined multiwavelet spaces. M2NO leverages multi-resolution analysis to selectively transfer low-frequency error components to coarser grids while preserving high-frequency details at finer levels. This design enhances both accuracy and computational efficiency without introducing additional complexity. Moreover, M2NO serves as an effective preconditioner for iterative solvers, further accelerating convergence in large-scale PDE simulations. Through extensive evaluations on diverse PDE benchmarks, including high-resolution, super-resolution tasks, and preconditioning settings, M2NO consistently outperforms existing models. Its ability to efficiently capture fine-scale variations and large-scale structures makes it a robust and versatile solution for complex PDE simulations. Our code and datasets are available on https://github.com/lizhihao2022/M2NO.
Partial differential equations, neural operator, multigrid, wavelets
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 9–13, 2025; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD 2026), August 9–13, 2025, Jeju Island, Republic of Koreaisbn: 979-8-4007-2258-5/2026/08doi: 10.1145/3770854.3780179ccs: Applied computing Physics
1. Introduction
Partial differential equations (PDEs) underpin models in fluid dynamics, solid mechanics, and climate science (82:intro; 87:turb; 20:turb; 02:pde). Classical solvers, like finite–difference and finite–element methods, offer provable convergence but can be prohibitively expensive for high-dimensional, multi-scale problems (12:fem). Multigrid algorithms mitigate this cost by eliminating low-frequency errors on successively coarser grids, often reducing the required Krylov iterations by orders of magnitude (00:tutorial; 77:bvp); however, they still rely on hand-crafted restriction/prolongation operators and demand problem-specific tuning for complex geometries. Recently, advancements in deep learning have introduced alternative approaches, such as Fourier Neural Operators (FNO) (21:fno), Galerkin Transformers (21:GT), and Deep Operator Networks (DeepONet) (21:DeepONet). These methods utilize neural network architectures to efficiently approximate mappings between parameters and solutions in infinite-dimensional spaces, tackling both forward and inverse problems in PDEs (19:pinn; 20:mgno; 23:no; 25:HarnessingSP).
Yet existing neural operators face three persistent hurdles: (i) their accuracy degrades when deployed on grids finer or coarser than those seen in training; (ii) they struggle to capture both global low-frequency trends and sharp local features within a single network; and (iii) they rarely match the iteration efficiency of hand-crafted multigrid solvers. To overcome these limitations, we propose an innovative framework that integrates a multigrid structure within predefined multiwavelet spaces, significantly enhancing the precision and efficiency of learning PDE operators.
Our methodology strategically transfers only the low-frequency error components to coarser grids for targeted elimination, drawing upon the robust principles of multigrid techniques (00:tutorial; 17:amg; 00:amg; 77:bvp). Traditional multigrid methods, effective yet often limited by scalability challenges in very large or complex scenarios, are enhanced by our proactive selection of appropriate coarse spaces and specific restriction/prolongation operators. This ensures that our multigrid process scales efficiently, optimizing both computational resources and adapting to complex problem sizes. In more detail, we employ a low-pass filter as the restriction operator, which specifically transfers the low-frequency components of errors into coarser grids, and use its transpose for the prolongation phase. This low-pass filter is ingeniously constructed using multiwavelets basis, providing a compact and data-efficient representation of operators across various scales (92:ten; 00:Fourier; 11:first; 21:mwt). The design process involves utilizing multiresolution analysis (MRA) to meticulously craft the multiwavelet transformation, integrating both the low-pass filter and its transpose as the fundamental restriction and prolongation operators (93:wave; 96:schur; 07:fastwave). This application of MRA within our framework significantly enhances our model’s ability to dynamically adapt to the evolving scales and complexities of PDE problems.
In this work, we introduce the Multiwavelet-based Multigrid Neural Operator (M2NO), a neural network architecture designed to inherently model multi-scale structures in PDEs, making notable contributions in the following areas:
  • Multiresolution and Multiscale Representation: We leverages MRA to design efficient restriction and prolongation operators based on multiwavelet transformations. By utilizing low-pass filters and their transposes, M2NO captures multi-scale information effectively, enabling precise mapping across different resolution levels. This structured representation allows for efficient error correction and improved adaptability in solving PDEs.
  • Enhanced Precision and Efficiency: M2NO significantly enhances the accuracy and efficiency of solving dynamic multiscale PDEs through the integration of a multigrid structure within predefined multiwavelet spaces. This strategic alignment enables precise error management by selectively addressing low-frequency components, thus optimizing the solution processes across varied scales and complexities.
  • Comprehensive Empirical Validation: M2NO has been rigorously evaluated across diverse 1D and 2D PDE benchmarks, including large-scale real-world datasets like ERA5, where it demonstrated strong performance in both multi-resolution and super-resolution tasks. Spectral Analysis further confirms M2NO’s effectiveness across different frequency bands, highlighting its robustness in capturing both low- and high-frequency components. Additionally, M2NO’s flexibility as a preconditioner for iterative solvers showcases its practical utility in accelerating convergence for complex PDE systems. These extensive experiments validate M2NO’s scalability, accuracy, and versatility in high-dimensional scientific computing.
2. Preliminaries
@@
In this section, we provide a concise overview of the multigrid method (17:amg; 00:tutorial) and multiresolution analysis using multiwavelets (02:mw; 92:ten), highlighting their similarities and how they inform the development of M2NO.
2.1. Multigrid Method
@@
The multigrid method is a well-grounded numerical technique for solving systems of equations that approximate PDEs. Specifically, we address the problem of solving the linear equation system:
(1) $A^{h}u^{h}=b^{h},$
@@
(3) $u^{h}=u^{h}+I_{2h}^{h}e^{2h}$
and perform a few more relaxation steps on the fine grid. This procedure outlines the foundational steps of the V-cycle multigrid scheme, which is recursively defined, as shown in Algorithm 1. For a comprehensive discussion and theoretical grounding of the Multigrid methods, please refer to the Appendix A.
2.2. Multiwavelets
@@
A MRA using multiwavelets involves a sequence of closed subspaces, defined as:
@@
(4) $\mathbf{V}_{n}^{k}\oplus\mathbf{W}_{n}^{k}=\mathbf{V}_{n+1}^{k},\quad\mathbf{W}_{n}^{k}\bot\mathbf{V}_{n}^{k},$
and the dimension of $\mathbf{W}_{n}^{k}$ is $2^{n}k$. Therefore, the composition of spaces is:
(5) $\mathbf{V}_{n}^{k}=\mathbf{V}_{0}^{k}\oplus\mathbf{W}_{0}^{k}\oplus\mathbf{W}_{1}^{k}\oplus\ldots\oplus\mathbf{W}_{n-1}^{k}.$
@@
In the two-dimensional discrete wavelet transform, depicted in Fig 5(a), $H$ and $G$ function as low-pass and high-pass filters, respectively, allowing the partitioning of frequency components at each level. Fig 5(b) visually represents this decomposition in a matrix format, enhancing our understanding of how submatrices $B_{n}$, $C_{n}$, and $D_{n}$ are derived from $L_{n}$.
The wavelet transform is defined as $\mathcal{W}:V_{j}\to V_{j-1}\oplus W_{j-1}$ (for $j\in\mathbb{Z}^{+}$), and is given by:
(6) $\mathcal{W}=\begin{pmatrix}H\\ G\end{pmatrix}$
@@
2.3. Similarities between Multigrid and MRA
@@
  1. (1)
    The highest resolution space $V_{n}$ in MRA aligns with $\Omega^{h}$, which represents the space of the finest grid vectors in the multigrid.
  2. (2)
    The decomposition in MRA, where data is broken down, corresponds to the restriction in multigrid. Similarly, the forward wavelet transform aligns with the restriction operator $I_{h}^{2h}$.
  3. (3)
    The reconstruction in MRA, which reassembles the data, is equivalent to the prolongation in multigrid. Correspondingly, the inverse wavelet transform is equivalent to the prolongation operator $I_{2h}^{h}$.
  4. (4)
    The orthogonality condition $\mathcal{W}\mathcal{W}^{T}=I$ corresponds to the identity relationship for the grid transfer operators $I_{h}^{2h}I_{2h}^{h}=I$.
  5. (5)
@@
These comparisons highlight the synergy between the multigrid and multiresolution properties of wavelet transforms and the error smoothing and correcting capabilities of intergrid operators in multigrid methods.
Figure 1. Illustration of M2NO with Two Levels and Grid Block. $f_{n}$ is initialized as $a_{n}$ and $u_{n}$ is initialized as 0. $R$ and $P$ are restriction and prolongation operator respectively. AA and $S$ are implemented as single-layer convolutional neural networks.
3. Methodology
This section explores the M2NO, leveraging the combined advantages of multigrid and multiresolution wavelet transformations. Initially, we define the problem and lay its theoretical groundwork in section 3.1. Section 3.2 explains the design and practical implementation of multiwavelet-based operators essential for M2NO. Subsequently, section 3.3 illustrates these methodologies’ application in solving discrete PDEs.
3.1. Problem Definition
@@
Formally, let $\mathcal{A}$ and $\mathcal{U}$ denote two Sobolev spaces $\mathcal{W}^{s,p}$ with $s>0$ and $p\geq 1$. In this context, the operator $\mathcal{G}$ maps from $\mathcal{A}$ to $\mathcal{U}$, $\mathcal{G}:\mathcal{A}\to\mathcal{U}$. We focus on cases where $s>0$ and $p=2$, choosing $p=2$ specifically to enable the definition of projections within a Hilbert space structure $\mathcal{H}^{s}$, using measures $\mu$.